### v2

# Out-of-Distribution Detection using Neural Activation Prior

Abstract
Out-of-distribution detection is a crucial technique for deploying machine learning models in the real world to handle the unseen scenarios. In this paper, we propose a simple but effective Neural Activation Prior (NAP) for out-of-distribution detection (OOD). Our neural activation prior is based on a key observation that, for a channel before the global pooling layer of a fully trained neural network, the probability of a few of its neurons being activated with a larger response by an in-distribution (ID) sample is significantly higher than that by an OOD sample. An intuitive explanation is each channel in a model fully trained on ID dataset would play a role in detecting a certain pattern in the samples within the ID dataset, and a few neurons can be activated with a large response when the pattern is detected in an input sample. Thus, a new scoring function based on this prior is proposed to highlight the role of these strongly activated neurons in OOD detection. This approach is plug-and-play and does not lead to any performance degradation on in-distribution data classification and requires no extra training or statistics from training or external datasets. Notice that previous methods primarily rely on post-global-pooling features of the neural networks, while the within-channel distribution information we leverage would be discarded by the global pooling operator. Consequently, our method is orthogonal to existing approaches and can be effectively combined with them in various applications. Experimental results show that our method achieves the state-of-the-art performance on CIFAR-10, CIFAR-100 and ImageNet datasets, which demonstrates the power of the proposed prior.
1 Introduction
Recently, deep learning has developed rapidly and become a crucial technique in various fields. However, neral networks frequently encounter data in inference that differs greatly from their training data, which is known as out-of-distribution (OOD) data. This challenge is growing more prevalent and is particularly vital in safety-critical areas such as autonomous driving [9, 16] and medical diagnosis [29], which urges the development of effective OOD detection methods.
Figure 1: Activation distribution difference between ID data and OOD data inside channel. This illustrates the difference in within-channel activation distribution between CIFAR-10 [19] (ID data) and the Texture [5] (OOD data) datasets. The horizontal axis represents the average activation value of the channel, while the vertical axis shows the maximal activation value of the channel. Additionally, semi-transparent shaded areas are introduced in the graph to represent the uncertainty surrounding each data point. As depicted, in this two-dimensional space formed by mean and maximum values, the ID and OOD data are linearly separable. Moreover, it is observed that, for identical mean values, the in-distribution data exhibit notably higher maximal activation values within the channel compared to the out-of-distribution data.
(a) Energy Score
(b) NAP Score
(c) Energy $\times$ NAP Score
Figure 2: Visualization of different score distributions obtained using Densenet [13] on ID dataset (CIFAR-10 [19]) and OOD datasets (iSun [40] and SVHN [27]) By simply multiplying (a) Energy Score with (b) NAP Score, the (c) Energy $\times$ NAP Score can achieve better results.This is because our proposed Neural Activation Prior is orthogonal to a series of existing OOD detection methods represented by Energy Score, so better results can be obtained by simple multiplication. It should be noted that our intention is not to compete with the Energy Score, but rather to emphasize that our proposed scoring function is complementary to it, as well as to other methods based on the Energy Score.
In the domain of out-of-distribution detection, OOD data refers to types of data that the model has not encountered during training. In practice, OOD data always has large diversity and is difficult to identify [41]. Researchers often phrase OOD detection as a one-class classification task, utilizing prior knowledge. Such methods [11, 23, 43] mainly rely on various priors to construct scoring functions to distinguish OOD samples from in-distribution samples. For example, Hendrycks et al. [11] observed that out-of-distribution samples typically exhibit lower maximum softmax probabilities, and accordingly proposed using the maximum softmax probability output by a neural network as an OOD indicator. Liu et al. [23] found that out-of-distribution samples usually have lower logits values, and based on this, an energy function was proposed for OOD detection. Although these methods have shown some promising results, there is still a gap in their application in real-world scenarios. At present, a mature and definitive paradigm for OOD detection has not yet established, encouraging us to explore various aspects of the OOD detection problem from different perspectives. Thus, in this paper, we aim to explore an effective OOD detection method to provide new insights into the field.
In this study, we propose a novel prior, called Neural Activation Prior (NAP), for out-of-distribution detection. Our neural activation prior is based on a key observation that for channels before the global pooling layer in a fully trained neural network (illustrated to Figure 3), a few neurons are significantly more likely to show a larger response when activated by in-distribution (ID) samples compared to OOD samples. An intuitive explanation is that channels in a model fully trained on an ID dataset play a role in detecting certain patterns in the input samples from the ID dataset. When such patterns are detected, a few neurons can be activated, resulting in larger responses. These large responses usually occurs when the input sample is ID data, but when the input is OOD data, such responses are rarely observed. This is because the pattern that the neuron is focusing on is unlikely to be present in the OOD data. To verify our proposed prior, we conduct experiments using the Densenet architecture [13] on the CIFAR-10 [19] and Texture [5] datasets. We compute the mean and maximum values of the within-channel activations before global pooling of all samples at the penultimate layer. As shown in Figure 1, it is obvious that the ID samples and OOD samples are linearly separable in the figure, and when the average activation value is the same, the maximal activation value of ID samples are significantly larger than that of OOD data.
It is worth noting that our proposed prior is orthogonal to that used in current OOD detection methods. In the field of out-of-distribution detection, as shown in Figure 3, existing methods [11, 22, 23, 34, 33, 8, 2] mainly focus on the outputs and weights of the neural network after global pooling, and use them to design scoring functions for OOD detection. In contrast, the prior we proposed is focus within the channels of the penultimate layer before global pooling. Since the information carried in our prior can be easily lost during the global pooling process, this demonstrates that our NAP is essentially complementary to the priors used in current OOD detection studies. This property is verified in subsequent experiments Section 4 in this paper.
Furthermore, in this paper, we propose a simple yet effective scoring function based on our prior NAP for OOD detection. To be precise, our scoring function is based on the ratio of the maximal and averaged activation values within a channel. It is a natural definition. On one hand, inspired by the concept of Signal-to-Noise Ratio (SNR), we can consider the maximal activation value as signal strength, while the averaged activation value represents noise strength. Therefore, the ratio of the maximum to the average value can be used to measure the quality of information contained in the channel. On the other hand, as shown in Figure 1, the ratio of the maximal and the averaged values for in-distribution samples is significantly higher than that of out-of-distribution data. Hence, from both the physical and practical standpoints, using this ratio as an OOD scoring function is a reasonable choice. It is noteworthy that this scoring function is a plug-and-play method, requiring no additional training, extra data, or reliance on pre-calculated statistical data from the training set. It does not lead to any degradation of the model’s inherent classification capabilities, and is complementary to existing methods and can be effectively combined with them to obtain better OOD detection results. For example, we can simply multiply our proposed scoring function with the energy function, and then we can significantly improve the OOD detection effect. Figure 2 shows a visualization of this result. These attributes greatly enhance the practicality of our approach, making our OOD detection technique applicable in a wide range of scenarios.
Experimental results show that our method achieves state-of-the-art performance on CIFAR-10 [19], CIFAR-100 [19] and ImageNet datasets. Specifically, our method significantly reduces the false positive rate by 40.06% (from 15.05 to 9.02) on the CIFAR-10 dataset [19], a reduction of 37.89% (from 41.40 to 25.71) on the CIFAR-100 dataset [19], and a reduction of 11.69% (from 35.66 to 31.49) on the ImageNet dataset. The large drop in FPR95 highlights the effectiveness of our approach in different environments. The above experimental results demonstrate the power of our proposed prior. We believe that these findings will provide inspiration for other researchers and thus promote progress in the field of OOD detection.
In summary, our contributions are as follows:
  • We propose a novel prior in OOD detection, named Neural Activation Prior (NAP). NAP is orthogonal to that used in existing OOD detection methods and we believe this prior can inspire researchers to further explore better OOD detection methods.
  • Based on the proposed prior, we develop a simple yet effective OOD detection scoring function. It is a plug-and-play approach that does not impair the model’s inherent capabilities. It can be readily integrated with many existing OOD detection techniques, enhancing their ability to balance OOD detection with in-distribution accuracy.
  • We demonstrate the state-of-the-art performance of our approach through extensive experiments across various datasets, highlighting its efficiency, ease of implementation, and effectiveness.
Discussion.
  • The scoring function proposed in this paper is a preliminary attempt to utilize the prior we proposed, aimed primarily at demonstrating its potential in OOD detection. There might be more effective scoring functions derived from this prior, but that is not the main concern of this paper.
  • We emphasize that the scoring function proposed in this paper is not intended to compete with functions proposed in existing work. Based on our proposed prior NAP, this scoring function is complementary to various existing works (such as a series of works [23, 34, 8] based on Energy Score). Therefore, our main focus is on integrating our approach with existing methods.
2 Related Work
2.1 Out-of-distribution detection
Research on out-of-distribution detection has explored a variety of techniques to underscore the distinctions between ID and OOD samples. These methods encompass classification-based [15, 22, 3, 7, 10], density-based [47, 1, 26, 46, 17, 28, 18, 31], and distance-based approaches [35, 21, 25, 30, 4, 14, 44, 37, 36, 24], with classification-based techniques generally outperforming the other types [41]. In classification-based methods, the basic work of OOD detection starts with a simple and effective baseline: using the Maximum Softmax Probability (MSP) [10] to measure the probability that a certain sample is an in-distribution sample. On this basis, early approaches [22, 12, 23] focused on developing enhanced OOD indicators derived from neural network outputs. In addition, some researchers have proposed strategies involving OOD sample generation [20] and gradient-based [22] techniques. Among these, certain post hoc methods [11, 22, 23, 34, 33, 8, 43, 2] are notable for their simplicity and because they do not necessitate changes in the training process or objectives. This feature is particularly valuable for implementing OOD detection in real production environments, where the additional cost and complexity associated with retraining would be unacceptable.
The Maximum Softmax Probability (MSP) method, initially presented by Hendrycks et al. [11], was a formative step in post hoc OOD detection, using a neural network’s softmax output as a heuristic for distinguishing ID from OOD samples. Its straightforward application facilitated early adoption in OOD studies. Despite MSP’s influence, its limitations prompted further innovation, giving rise to the Energy method. This method, proposed by Liu et al. [23], refines the approach by assigning an energy score to network outputs, showing quantitative improvements over MSP with theoretical and empirical support. Advancements in post hoc OOD detection have led to diverse methodological branches stemming from MSP [10] and Energy [23] paradigms. LINe [2] innovates by reducing neuron-induced noise through the calculation of Shapley values. Yu et al. [43] distinguish ID from OOD data by identifying neural network blocks with optimal differentiation based on the norms of their features. DICE [33] improves discrimination by pruning weights in the fully connected layer according to the contribution units make during classification. On the other end of the spectrum, entirely computation-free post hoc methods such as ReAct [34] and ASH [8] have shown promise. ReAct [34] investigates activations prior to the fully connected layer, applying rectification to suppress extreme activations that OOD data tend to trigger, thereby achieving refined detection outcomes. Similarly, ASH [8] prunes the activations inputted to the fully connected layer, but it achieves even more enhanced results compared to DICE [33] by its selective pruning strategy. In this paper, our comparison mainly focuses on post hoc methods, since our method also belongs to this category.
3 Neural Activation Prior
The neural activation prior is based on the following observation for OOD detection: for a channel located before the global pooling layer in a fully trained neural network, the likelihood that a small number of its neurons activated with a stronger response to an in-distribution (ID) sample is significantly higher compared to an out-of-distribution (OOD) sample. For the behavior in other layers of the neural network, refer to the discussion around Figure 5 in Section 5.5.2.
To formally describe this observation, we first define the concept of neural activation. Consider a trained classification neural network $f$, assuming it receives $D$-dimensional input $x$ and outputs $K$-dimensional logits. That is $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{K}$. We concentrate on the activation tensor $\mathbf{A}$ , located at the penultimate layer just before the global pooling operation, as illustrated in Figure 3. Let the dimensions of $\mathbf{A}$ be $C\times H\times W$, where $C$ is the number of channels, and $H$ and $W$ are the spatial dimensions.
Figure 3: Illustration identifying the focus zone of Neural Activation Prior (NAP) in a classification neural network. The figure highlights the specific position of NAP in the network. The activation value in the green part is the location of the NAP proposed in this article, which is before global pooling. In contrast, most existing methods primarily focus on the regions after global pooling.
Let $\mathbf{A}_{j}$ represent the activation tensor of the $j$-th channel. We define two key statistical indicators of $\mathbf{A}_{j}$ as follows:
  • Maximum Activation Value:
    $\text{Max}(\mathbf{A}_{j})=\max_{k,l}\mathbf{A}_{jkl},$
    where $\text{Max}(\mathbf{A}_{j})$ is the maximum value among all elements in the activation vector $\mathbf{A}_{j}$. Notice that $\mathbf{A}_{jkl}$ is processed by ReLU (as shown in Figure 3), hence $\mathbf{A}_{jkl}\geq 0$.
  • Mean Activation Value:
    $\text{Mean}(\mathbf{A}_{j})=\frac{1}{h\times w}\sum_{k=1}^{h}\sum_{l=1}^{w}\mathbf{A}_{jkl},$
    where $\text{Mean}(\mathbf{A}_{j})$ is the average of all activation values in the $j$-th channel.
Inspired by the concept of signal-to-noise ratio, we interpret $\text{Max}(\mathbf{A}_{j})$ as the signal strength and $\text{Mean}(\mathbf{A}_{j})$ as the noise strength. The ratio of these values can be viewed as an indicator of the quality of the activation vector $\mathbf{A}_{j}$. Note that we calculate this metric separately for each channel in the network, since different channels are usually used to detect different objects. In the context of OOD detection, this quality measure can be used to assess whether the neural network recognizes the input sample — in other words, this measure can be used to judge whether a sample is within the training set distribution. As shown in Figure 1, this phenomenon is termed as the neural activation prior.
It is worth noting that our proposed prior is orthogonal to existing OOD detection methods. As illustrated in Figure 3, existing methods mainly focus on the network output and weight of the penultimate layer after global pooling, and leverage them to design various scoring functions for OOD detection. In contrast, the prior we propose focuses on the channels of the penultimate layer before global pooling. Since the distribution information within these channels is inevitably lost during the global pooling process, our proposed prior is complementary to existing work.
4 OOD Detection with NAP
4.1 Basics
First, we will provide a brief overview of typical settings for out-of-distribution (OOD) detection in image classification networks. Typically, classification networks use in-distribution (ID) data, that is, known training data sets, to train a classification model. Once training is completed, the model’s parameters are fixed, enabling it to effectively classify the categories within the training data set. In the testing phase, to identify OOD samples, researchers usually introduce a scoring function into the model. During the inference, samples mixed with OOD data are fed into this trained model. The model not only classifies each sample, but also uses a scoring function to generate a score for each one. This score is used to evaluate whether a sample belongs to an in-distribution class in the training set, or an unknown out-of-distribution class.
4.2 Design of scoring function
Based on our prior proposed previously in Section 3, we propose a SNR-like scoring function. In our formulation, the mean activation value is interpreted as the noise intensity, while the maximal activation value is regarded as the signal strength. This conceptual framework leads to the following scoring function:
[…]
where $C$ represents the number of channels before global pooling. In this formula, a small constant $\epsilon>0$ is added to ensure the numerical stability of the computation.
[…]
We can combine these two scoring functions. In this paper, we adopt the weighted geometric mean method to combine them:
[…]
Discussion.
  • Plug-and-Play Simplicity: The scoring function we proposed is a plug-and-play approach that can be easily integrated into existing neural network architectures. It requires no additional training or external data and retains the model’s inherent classification capabilities. These properties make it practical and suitable for a variety of applications.
  • Orthogonal to Existing Approaches: Based on our proposed priors, the scoring function we design is orthogonal to existing methods. As shown in Figure 1, the value ranges of the within-channel activation mean values of ID samples and OOD samples overlap. This puts existing methods into trouble when distinguishing between ID and OOD samples with close means values. Based on the prior we proposed, this kind of dilemma can be solved naturally, which illustrates the power of our proposed priors to provide new perspectives for identifying OOD data.
5 Experiments
Table 1: Comparison with competitive post hoc out-of-distribution detection method on CIFAR [19] benchmarks. All values are percentages and are averaged over 6 OOD test datasets. Except for our proposed method, all other results are taken from [8]. The best result is highlighted in bold. Methods marked with an asterisk (*) are employed with the Energy Score [23].
Method CIFAR-10 CIFAR-100
FPR95 AUROC FPR95 AUROC
MSP [11] 48.73 92.46 80.13 74.36
ODIN [22] 24.57 93.71 58.14 84.49
Mahalanobis [21] 31.42 89.15 55.37 82.73
Energy [23] 26.55 94.57 68.45 81.19
ReAct* [34] 26.45 94.67 62.27 84.47
DICE* [33] 20.83 95.24 49.72 87.23
ASH-B* [8] 20.23 96.02 48.73 88.04
ASH-S* [8] 15.05 96.91 41.40 90.02
NAP-E (Ours) 9.02 98.15 32.61 92.84
NAP-A (Ours) 11.14 97.48 35.40 91.21
NAP-R (Ours) 9.18 98.02 25.71 93.18
Table 2: OOD detection results on ImageNet-1k [6]. All values are percentages. Except for our proposed method, all other results are taken from [8]. The best result is highlighted in bold. Methods marked with an asterisk (*) are employed with the Energy Score [23].
Method OOD Datasets Average
iNaturalist [38] SUN [39] Places [45] Textures [5]
FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC
MSP [11] 64.29 85.32 77.02 77.10 79.23 76.27 73.51 77.30 73.51 79.00
ODIN [22] 55.39 87.62 54.07 85.88 57.36 84.71 49.96 85.03 54.20 85.81
Mahalanobis [21] 62.11 81.00 47.82 86.33 52.09 83.63 92.38 33.06 63.60 71.01
Energy [23] 59.50 88.91 62.65 84.50 69.37 81.19 58.05 85.03 62.39 84.91
ReAct* [34] 42.40 91.53 47.69 88.16 51.56 86.64 38.42 91.53 45.02 89.47
DICE* [33] 43.09 90.83 38.69 90.46 53.11 85.81 32.80 91.30 41.92 89.60
ASH-B* [8] 31.46 94.28 38.45 91.61 51.80 87.56 20.92 95.07 35.66 92.13
ASH-S* [8] 39.10 91.94 43.62 90.02 58.84 84.73 13.12 97.10 38.67 90.95
NAP-E (Ours) 29.90 94.47 39.69 90.46 55.17 85.15 11.74 97.28 34.12 91.84
NAP-A (Ours) 28.34 94.69 36.29 91.97 51.25 87.27 12.36 97.10 32.06 92.76
NAP-R (Ours) 24.58 95.55 38.47 91.12 53.32 86.24 9.57 97.60 31.49 92.63
5.1 Evaluation on CIFAR benchmarks
Implementation Details. In our experiments, consistent with recent studies [34, 33, 8], we utilized 10,000 test images from both CIFAR-10 [19] and CIFAR-100 [19] as in-distribution data. To gauge the performance of the model, six widely-used OOD datasets were employed as benchmarks. These datasets include SVHN [27], Textures [5], iSUN [40], LSUN-Crop [42], LSUN-Resize [42], and Places365 [45]. As for pre-trained model, we employed DenseNet [13], and we follow the training setting of DenseNet introduced in [33].
[…]
Comparison. For comparison, we adopted recent post-hoc OOD detection methods: MSP [11], ODIN [22], Mahalanobis distance [21], Energy [23], ReAct [34], DICE [33], and ASH [8].
Experimental Results. Table 1 shows the comparison of NAP with other post hoc OOD detection methods on the CIFAR-10 and CIFAR-100 benchmarks. As shown in the table, our method significantly outperforms all other methods on the CIFAR-10 and CIFAR-100 datasets, achieving state-of-the-art performance. In CIFAR-10, NAP-E reduces FPR95 by 40.06% (from 15.05 to 9.02) compared to the previous best method. In CIFAR-100, NAP-R reduces FPR95 by 37.89% (from 41.40 to 25.71) compared to the previous best method. What is impressive is that the implementation of NAP-E is extremely simple and does not cause any degradation in ID data classification accuracy. This enables NAP-E to achieve the optimal pareto front among existing methods, where relevant results can be seen in Section 5.3. The Above results demonstrate the effectiveness of our method and the power of NAP in OOD detection.
5.2 Evaluation on ImageNet
Implementation Details. In real-world applications, models are confronted with high-resolution images spanning a diverse range of scenes and features. Evaluations on large-scale datasets can provide insights into the performance of models in practical deployments. Thus, in line with recent research [34, 33, 8], we conduct experiments with NAP on the expansive ImageNet-1k [6] dataset in this study. Four dataset subsets, with all overlapping categories with ImageNet-1k eliminated, were employed as OOD benchmarks. These OOD datasets comprise Textures [5], Places365 [45], iNaturalist [38], and SUN [39]. We used MobileNetV2 [32] architecture, which pre-trained on ImageNet-1k [6]. The architecture and parameters remain unchanged during the OOD detection stage. Parallel to the experiments conducted on the CIFAR [19] benchmarks, we experimented with three distinct implementations. For parameter settings, $w$ was set to $0.6$ for NAP-E, $0.8$ for NAP-A, and $0.8$ for NAP-R.
Experimental Results. Table 2 shows the comparison results of NAP with other post hoc OOD detection methods on the ImageNet-1k [6] benchmark. As shown in the table, our method outperforms all other methods on average on the ImageNet-1k dataset [6], achieving state-of-the-art performance. We find that our method is less effective on the Places dataset [45]. We argue that this is because the scene in Places [45] is complex and contains some patterns that can cause large response values. These patterns are close to the patterns that the original network focuses on, so it is relatively difficult to distinguish them.
(a) CIFAR-10
(b) CIFAR-100
Figure 4: Investigating the trade-offs between In-Distribution (ID) classification accuracy and Out-Of-Distribution (OOD) detection rate on CIFAR benchmarks [19] across various methods. All methods and experiments are implemented by us. Methods prefixed with ’NAP’ are visually distinguished, highlighted in various shades of red in the figures.
We also note that all three of our implementations outperform the state-of-the-art methods on the iNaturalist [38] and Textures [5] datasets. Intuitively, most of the pictures in the Textures [5] dataset are of a texture nature, so there is a small probability of obtaining a large response value. And while samples in iNaturalist have a relatively simple background, and the animals and plants in the foreground have certain semantic differences from the samples in ImageNet-1k, therefor, a sample in iNaturalist [38] is not likely to trigger large response values. The reason why existing methods do not perform well on this dataset may be due to their focus on activation values after global pooling. We speculate that although texture pictures cannot cause large activation values, they can excite small-amplitude noise on the entire feature map. Since large response values tend to only appear in a small area on the feature map, after global pooling, in-distribution samples and texture class samples are likely to obtain similar average activation values. This may compromise the separability of the two types of samples, causing existing work to perform poorly on such samples.
5.3 Evaluation on Pareto Frontier of ID accuracy and OOD detection
The goal of OOD detection methods is to provide the best OOD detection performance without compromising the inherent classification performance of the original model. While some existing methods improve OOD detection performance, they also impair the model’s classification ability within the distribution. Although the authors of these methods claimed in their papers that the damage was minor, when we tried to combine different methods, the subtle changes to the model by different methods eventually led to a significant decrease in the classification ability of the original model. This cumulative damage hinders the possibility of combining multiple methods for OOD detection. Based on NAP, our proposed method preserves the model’s classification ability and is simple to implement, thereby offering a cost-free enhancement in OOD detection performance. Figure 3(a) and Figure 3(b) highlight the trade-off between in-distribution (ID) classification accuracy and out-of-distribution (OOD) detection rate (AUROC) on the CIFAR benchmark [19]. Our proposed NAP family demonstrates the optimal balance of this trade-off, demonstrating superior performance along the Pareto front.
[…]
Table 3: Comparison with competitive out-of-distribution score method on CIFAR benchmarks [19] and ImageNet-1k [6] dataset. CIFAR results are averaged across 6 different OOD tasks and ImageNet-1k results are averaged across 4 different OOD tasks.
Method CIFAR-10 CIFAR-100 ImageNet-1k
FPR95 AUROC FPR95 AUROC FPR95 AUROC
MSP [11] 48.69 92.52 80.06 74.45 73.51 79.00
Energy [23] 26.59 94.63 68.29 81.23 62.39 84.91
[…]
(a) After first convolution layer
(b) In Transaction Block 1
(c) In Transaction Block 2
(d) Before Global Pooling
Figure 5: Activation distribution at different positions within the DenseNet architecture [13] applied to CIFAR-10 [19] and Places365 datasets [45]. For this analysis, four specific locations within the network were chosen: (a) after the first convolution layer, (b) just before the pooling operation in the first transition block, (c) just before the pooling operation in the second transition block, and (d) right before the final global pooling layer. It is observed that with increasing depth, the separability between in-distribution and out-of-distribution samples becomes more pronounced in the two-dimensional space defined by the maximum and average activation values.
5.5 Ablation studies
5.5.1 Ablation on different parameter $w$.
Table 4: Ablation results on different $w$ of NAP-R on CIFAR [19] benchmarks and ImageNet-1k [6] dataset. CIFAR results are averaged across 6 different OOD tasks and ImageNet-1k results are averaged across 4 different OOD tasks.
$w$ CIFAR-10 CIFAR-100 ImageNet-1k
FPR95 AUROC FPR95 AUROC FPR95 AUROC
0.0 26.57 92.45 54.91 85.86 49.91 83.07
0.1 20.36 94.77 49.80 87.75 48.59 84.21
0.2 14.61 96.49 44.03 89.58 46.82 85.49
0.3 10.83 97.53 36.81 91.21 44.88 86.92
0.4 9.18 98.02 29.66 92.47 42.43 88.44
0.5 9.47 98.11 25.71 93.18 39.42 89.96
0.6 10.72 97.93 26.83 93.15 36.02 91.33
0.7 12.98 97.53 32.02 92.13 33.03 92.32
0.8 16.92 96.91 43.99 89.86 31.49 92.63
0.9 21.89 96.04 57.95 86.37 34.38 91.78
1.0 29.00 94.92 69.94 82.07 49.90 88.75
In Table 4, we show the various OOD detection performances of NAP-R on CIFAR10 [19], CIFAR100 [19] and ImagetNet-1k [6] by varying the threshold $w$. Experimental results show that by choosing an appropriate $w$, we can obtain better OOD detection results by combining NAP with the existing OOD scoring function. This intuitively indicates that our proposed prior is orthogonal to and effectively complements existing methods. See the Appendix for the performance impact of changes in $w$ for other implementations.
5.5.2 Does NAP Work across all network layers?
In our study, we visualized the distribution of activations at different positions in the DenseNet architecture applied to the CIFAR-10 [19] and Places365 datasets [45]. Four specific locations were chosen for this analysis: after the first convolutional layer, before the pooling operation in the first transition block, before pooling in the second transition block, and before the final global pooling layer. We selected some typical cases from the experimental results to display in Figure 5. For a more comprehensive set of visualization results, please see the Appendix. From Figure 5, we can observe that as the position becomes deeper, the separability of in-distribution samples and out-of-distribution samples becomes more and more obvious in the two-dimensional space spanned by the maximum and average activation values. This is because in the first few layers of the network, neurons tend to capture some low-level features, and such features are likely to be present in both ID and OOD samples. Therefore, as seen in Figures 4(a) and 4(b), the two lines intersect. As we examine deeper neurons in the network, the features captured by neurons become features containing certain semantic information. Therefore, as shown in  4(c) and 4(d), data within the distribution with specific semantic information can obtain relatively larger response values. This suggests that our proposed Neural Activation Prior is more effective towards the end of the network, particularly before the global pooling layer.
6 Conclusion
This paper proposes a novel Neural Activation Prior (NAP) for out-of-distribution (OOD) detection in machine learning models. Our proposed prior is grounded on the observation that in a fully trained neural network, in-distribution (ID) samples typically induce stronger activation responses in some neurons of a channel compared to out-of-distribution samples. This discovery led to our novel scoring function based on within-channel distribution. Its main advantage lies in its simplicity and easy integration. It necessitates neither additional training nor external data and does not compromise the classification performance on ID data. Experimental results on various datasets and architectures show that our method achieves state-of-the-art performance in OOD detection. This not only verifies the effectiveness of neural activation priors, but also demonstrates the potential of rethinking the way neural network features are utilized in OOD scenarios.
Supplementary Material
Table 5: Scoring function proposed based on NAP is compatible with and improves on existing methods on CIFAR-10 dataset. All methods and experiments are implemented by us. All values in this table are percentages. The average over six OOD test datasets is also reported. The methods with prefix ’NAP-’ (e.g., NAP-E, NAP-A, NAP-R) represent the integration of NAP with various existing methods (Energy [23], ASH-S [8], ReAct [34] respectively).
Method OOD Datasets Average
SVHN Textures iSUN LSUN LSUN-Crop Places365
FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC
Energy 40.57 93.99 56.29 86.42 10.07 98.07 9.28 98.12 3.81 99.15 39.50 92.01 26.59 94.63
NAP-E (Ours) 8.32 98.36 11.65 97.72 1.77 99.57 1.50 99.60 0.99 99.76 29.89 93.91 9.02 98.15
ASH-S 6.51 98.65 24.34 95.09 5.17 98.90 4.96 98.92 0.90 99.73 48.45 88.34 15.05 96.91
NAP-A (Ours) 5.55 98.86 10.51 97.90 3.04 99.32 2.68 99.40 0.80 99.80 44.28 89.59 11.14 97.48
ReAct 41.64 93.87 43.58 92.47 12.72 97.72 11.46 97.87 5.96 98.84 43.31 91.03 26.45 94.67
NAP-R (Ours) 8.07 98.31 8.10 98.17 2.81 99.35 2.35 99.43 3.04 99.33 30.70 93.50 9.18 98.02
Table 6: Scoring function proposed based on NAP is compatible with and improves on existing methods on CIFAR-100 dataset. All methods and experiments are implemented by us. All values in this table are percentages. The average over six OOD test datasets is also reported. The methods with prefix ’NAP-’ (e.g., NAP-E, NAP-A, NAP-R) represent the integration of NAP with various existing methods (Energy [23], ASH-S [8], ReAct [34] respectively).
Method OOD Datasets Average
SVHN Textures iSUN LSUN LSUN-Crop Places365
FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC
Energy 87.46 81.85 84.15 71.03 74.54 78.95 70.65 80.14 14.72 97.43 79.20 77.72 68.45 81.19
NAP-E (Ours) 19.03 96.40 21.72 95.47 33.24 94.15 43.38 92.11 2.60 99.38 75.70 79.54 32.61 92.84
ASH-S 25.02 95.76 34.02 92.35 46.67 91.30 51.33 90.12 5.52 98.84 85.86 71.62 41.40 90.02
NAP-A (Ours) 17.41 96.72 22.70 94.99 38.22 93.34 43.05 92.17 5.25 98.94 85.76 72.08 35.40 91.32
ReAct 83.81 81.41 77.78 78.95 65.27 86.55 60.08 87.88 25.55 94.92 82.65 74.04 62.27 84.47
NAP-R (Ours) 14.19 96.52 17.22 96.16 16.72 96.54 17.16 96.64 5.73 98.76 82.54 74.46 25.71 93.18
In this appendix, we provide comprehensive additional materials to supplement the main text. The contents include:
  • Pseudo Code for Neural Activation Prior (Section A): The algorithmic representation of the Neural Activation Prior (NAP) methodology.
  • Enhancing Existing Methods with NAP: A CIFAR Benchmark Case Study (Section B): A case study demonstrating the application of NAP to enhance existing methods, with a focus on the CIFAR benchmark.
  • Why Is the Penultimate Layer More Effective for Neural Activation Prior? (Section C): An exploration of the reasons behind the superior effectiveness of the penultimate layer in the context of NAP.
  • Evaluating Multi-Layer Integration with NAP for OOD Detection (Section D): Investigation into the effects of integrating multiple layers along with NAP in detecting OOD data.
  • On Transferability to Other Architectures (Section E): Discussion on how the NAP approach can be adapted to different neural network architectures.
  • Full Results of Ablation on Parameter $w$ (Section F): Detailed analysis and results concerning the impact of varying the parameter $w$ on our model’s performance.
  • Discussion (Section G): A concluding section that summarizes the key findings and outlines future directions for research based on our work.
Appendix A Pseudo Code for Neural Activation Prior
As illustrated in Algorithm 1, we present a detailed pseudo-code representation of our proposed method for OOD detection, which is integrated into the DenseNet architecture. The key modification involves the calculation of NAP score $\mathcal{S}$ within the DenseNet’s processing pipeline (highlighted in green font in the algorithm), which is then followed by calculating the OOD score using the model’s logits together with $\mathcal{S}$. These calculations do not alter the logits output by the model, thereby ensuring no degradation in classification accuracy for the ID dataset.
Algorithm 1 OOD Detection Using Neural Activation Prior on DenseNet
0:   Image $x$, Weight $w$ (for OOD score calculation)
0:   Output logits, OOD Score
1:   Apply initial layers of DenseNet on $x$ to obtain intermediate output:
2:      $out\leftarrow\text{conv1}(x)$
3:      $out\leftarrow\text{trans1}(\text{block1}(out))$
4:      $out\leftarrow\text{trans2}(\text{block2}(out))$
5:      $out\leftarrow\text{block3}(out)$
6:      $out\leftarrow\text{relu}(\text{bn1}(out))$
7:      Let $\mathcal{A}$ denote this intermediate output.
8:   Compute NAP score $\mathcal{S}$ from $\mathcal{A}$:
9:      Flatten $\mathcal{A}$ across dimensions 2 and 3.
10:      Compute Max of $\mathcal{A}$ across flattened dimensions.
11:      Compute Mean of $\mathcal{A}$ over dims 2, 3.
12:      Calculate $\mathcal{S}$ as $(\text{Max of }\mathcal{A}/(\text{Mean of }\mathcal{A}+1))^{2}$.
13:      Compute the mean of $\mathcal{S}$ across dimension 1.
14:   Continue with DenseNet forward pass:
15:      Apply average pooling and reshape on $\mathcal{A}$.
16:      Get logits from fully connected layer.
17:   Calculate OOD Score:
18:      Compute log-sum-exp of the logits.
[…]
20:   return  Output logits, OOD Score
Appendix B Enhancing Existing Methods with NAP: A CIFAR Benchmark Case Study
This section focuses on the significant enhancements brought by the Neural Activation Prior (NAP) to existing out-of-distribution detection methods in the context of the CIFAR-10 (in Table 5) and CIFAR-100 (in Table 6) dataset. The incorporation of NAP into established methods like Energy [23], ASH-S [8], and ReAct [34], resulting in NAP-E, NAP-A, and NAP-R respectively, showcases the potential of NAP in augmenting existing approaches. Our experimental results demonstrate that NAP-based variants consistently outperform their corresponding traditional methods across all six OOD datasets. Notably, our experimental results reveal some extremely substantial decreases in FPR95 values, indicative of the profound impact of NAP integration. For instance, on the CIFAR-100 dataset, the FPR95 value of NAP-R, compared to ReAct, dropped by 83.06% (from 83.81 to 14.19), highlighting a notable reduction in false alarms while maintaining high detection accuracy and affirming the enhanced capability of these methods in distinguishing between in-distribution and OOD samples.
(a) After first convolution layer
(b) In Transaction Block 1
(c) In Transaction Block 2
(d) Before Global Pooling
Figure 6: Activation distribution at different positions within the DenseNet architecture [13] applied to CIFAR-10 [19] and Places365 datasets [45]. For this analysis, four specific locations within the network were chosen: (a) after the first convolution layer, (b) just before the pooling operation in the first transition block, (c) just before the pooling operation in the second transition block, and (d) right before the final global pooling layer. Note that we only include data points with an average activation over $0.1$. As shown in the figure, the first three selected layers show a less marked distinction between ID and OOD samples, while the fourth layer, preceding the final global pooling layer, demonstrates clearer separability and enhanced stability. Therefore, the fourth selected layer (the penultimate layer) is more suitable for developing a scoring function for OOD detection.
Appendix C Why Is the Penultimate Layer More Effective for Neural Activation Prior?
We provide an extensive collection of visualizations showcasing the activations within the DenseNet architecture when applied to the CIFAR-10 (ID) dataset and Places365 (OOD) dataset. These visualizations are crucial for understanding how the network processes both ID and OOD data, revealing the distinct patterns of neural activations at various layers of the network. Our analysis focuses on four critical layers within DenseNet: (1) after the first convolutional layer, (2) before the pooling operation in the first transition block, (3) before pooling in the second transition block, and (4) before the final global pooling layer. Each layer offers four visualizations, providing a comprehensive view of the network’s response to different datasets.
Figure 7: Activation distribution at penultimate layer before global pooling operation within the MobileNetV2 architecture [32] applied to ImageNet-1k and iNaturalist datasets [38]. We only include data points with an average activation over $0.1$. The figures show that our Neural Activation Prior (NAP) method is also effective in MobileNetV2, proving that NAP can be applied to different architectures.
These detailed visualizations enhance the discussion in the main text, offering deeper insights into how NAP effectively distinguishes between ID and OOD samples within the network. As depicted in Figure 6, the first three selected layers, which focus primarily on low-level features, exhibit a less pronounced distinction between ID and OOD samples. This is likely because low-level features, such as edges and textures, are common to both ID and OOD datasets, making them less distinctive. However, the contrast between ID and OOD samples becomes more evident and stable in the fourth selected layer, located before the final global pooling layer. This layer, concentrating on high-level semantic information, captures features more unique to the ID dataset, leading to clearer separability and enhanced stability in activation values compared to earlier layers. This layer’s focus on distinctive semantic features makes it particularly suitable for developing a scoring function for OOD detection. Therefore, since the penultimate layer is the most informative layer in the neural network, we utilize this layer in our method to develop our scoring function.
Table 7: Experimental results of OOD detection using NAP at different layers in DenseNet [13] on CIFAR-10 and CIFAR-100 datasets. NAP(x) indicates the computation of OOD score at the layer ’x’, where ’c1’ corresponds to after the first convolutional layer, ’t1’ before the pooling operation in the first transition block, ’t2’ before pooling in the second transition block, and ’p’ before the final global pooling layer. Combinations of layers, indicated by commas in NAP(..), represent the multiplication of OOD scores from respective layers. These selected layers are consistent with those used for visualizations described earlier in the paper. Notably, NAP(p) is the approach actually utilized in our paper.
Method CIFAR-10 CIFAR-100
FPR95 AUROC FPR95 AUROC
NAP(c1) 83.22 51.99 84.13 50.34
NAP(t1) 69.10 50.47 86.82 54.92
NAP(t2) 56.53 78.44 88.85 53.08
NAP(c1,t1,t2,p) 68.33 58.99 82.66 56.96
NAP(t1,t2,p) 56.84 64.26 83.35 57.97
NAP(t2,p) 34.41 87.81 82.74 58.31
NAP(p) 26.57 92.45 54.91 85.86
Appendix D Evaluating Multi-Layer Integration with NAP for OOD Detection
In an additional exploration presented in the appendix, we investigate the effects of incorporating values from both the penultimate layer and earlier layers on OOD detection. Our experiments, detailed in Table 7, suggest that the integration of earlier layers with the penultimate layer, where NAP is primarily applied, does not yield significant improvements in OOD detection. This phenomenon could partly stem from the inherent limitations of earlier layers in differentiating between ID and OOD data. Additionally, there is a possibility that the scoring function, specifically optimized for the penultimate layer, may not align optimally with the feature representation characteristics of the preceding layers. Considering the constraints of space in this paper, a comprehensive analysis of multi-layer integration using NAP is not presented. Nonetheless, the potential of combining multiple layers in OOD detection, especially in the context of NAP, remains an intriguing aspect for future research. We anticipate that further investigations, potentially involving the creation of new scoring functions suitable for a broader range of layers, could provide substantial contributions to the field. Thus, we propose this as an avenue for future work, aiming to stimulate further advancements within the research community.
Appendix E On Transferability to Other Architectures
In Figure 7, 8, and 9, we present detailed visualizations of the activation patterns within three distinct architectures: MobileNetV2 [32], ResNet50 [ r1], and VGG16 [ r2]. These visualizations clearly demonstrate a remarkable gap between ID samples, depicted with blue lines, and OOD samples, represented with orange lines. This distinction is evident across all three architectures, underscoring the versatility and effectiveness of the proposed Neural Activation Prior (NAP). The consistent separability observed in these diverse architectures confirms the adaptability and potential of NAP for broad application in different neural network models.
Table 8: Ablation results on different $w$ of NAP-E on CIFAR [19] benchmarks and ImageNet-1k dataset. CIFAR-10 [19] and CIFAR-100 [19] results are averaged across 6 different OOD tasks and ImageNet-1k results are averaged across 4 different OOD tasks. The rows with $w=0.0$ and $w=1.0$ are actually the results of NAP and Energy [23], respectively. The results in bold demonstrate that properly combing NAP and Energy [23] can lead to significant better results, which implies that orthogonality between NAP and the existing methods.
$w$ CIFAR-10 CIFAR-100 ImageNet-1k
FPR95 AUROC FPR95 AUROC FPR95 AUROC
0.0 26.57 92.45 54.91 85.86 49.91 83.07
0.1 17.94 95.44 47.95 88.85 47.35 85.01
0.2 11.64 97.19 40.24 91.14 44.39 86.99
0.3 9.26 97.96 34.24 92.50 41.18 88.84
0.4 9.02 98.15 32.61 92.84 37.99 90.37
0.5 9.71 98.00 35.13 92.19 35.25 91.41
0.6 11.85 97.63 42.30 90.63 34.12 91.84
0.7 14.56 97.08 49.50 88.48 35.50 91.61
0.8 19.20 96.37 57.11 86.05 40.01 90.69
0.9 22.31 95.55 63.21 83.59 48.05 89.02
1.0 26.58 94.63 68.29 81.23 58.87 86.60
Appendix F Full Results of Ablation on Parameter $w$
In Table 8 and Table 9, we show the various OOD detection performances of NAP-E and NAP-A on CIFAR10 [19], CIFAR100 [19] and ImagetNet-1k by varying the threshold $w$. Experimental results demonstrate that by choosing an appropriate $w$ to combine NAP with existing OOD scoring function, we can obtain significant better results than both NAP (i.e., $w=0.0$ in the table) and existing OOD score based method (i.e., $w=1.0$ in the table). This intuitively indicates that our proposed prior is orthogonal to and effectively complements existing methods.
Table 9: Ablation results on different $w$ of NAP-A on CIFAR [19] benchmarks and ImageNet-1k dataset. CIFAR-10 [19] and CIFAR-100 [19] results are averaged across 6 different OOD tasks and ImageNet-1k results are averaged across 4 different OOD tasks. The rows with $w=0.0$ and $w=1.0$ are actually the results of NAP and ASH [8], respectively. The results in bold to significant better results, which implies that orthogonality between NAP and the existing methods.
$w$ CIFAR-10 CIFAR-100 ImageNet-1k
FPR95 AUROC FPR95 AUROC FPR95 AUROC
0.0 26.57 92.45 54.91 85.86 49.91 83.07
0.1 20.21 94.59 49.45 87.90 47.44 85.15
0.2 15.60 96.02 44.48 89.36 44.62 87.14
0.3 12.99 96.87 40.55 90.32 41.55 88.90
0.4 11.60 97.31 37.43 90.89 38.66 90.36
0.5 11.14 97.48 36.13 91.16 35.83 91.47
0.6 11.32 97.47 35.40 91.21 33.50 92.22
0.7 11.98 97.35 35.66 91.08 32.28 92.64
0.8 12.91 97.15 37.43 90.82 32.06 92.76
0.9 14.03 96.90 39.44 90.45 33.13 92.58
1.0 15.06 96.61 41.43 90.01 35.74 92.13
Figure 8: Activation distribution at penultimate layer before global pooling operation within the ResNet50 architecture applied to ImageNet-1k and iNaturalist datasets [38]. We only include data points with an average activation over $0.1$. The figures show that our Neural Activation Prior (NAP) method is also effective in ResNet50, proving that NAP can be applied to different architectures.
Figure 9: Activation distribution at penultimate layer before global pooling operation within the VGG16 architecture applied to ImageNet-1k and iNaturalist datasets [38]. We only include data points with an average activation over $0.1$. The figures show that our Neural Activation Prior (NAP) method is also effective in VGG16, proving that NAP can be applied to different architectures.
Appendix G Discussion
Based on the ultra effectiveness of our prior, shown in the visualization results presented in Figure 5 in the main text and Figure 7, 8, and 9 in this appendix, we are confident that more effective scoring functions exist. Due to space limitations, the primary focus and contribution of this paper is to introduce this prior and and validate its effectiveness by proposing a new score function. We leave the development of optimal scoring functions based on our prior for future work, aiming to further contribute to the community. We also hope this paper could encourage the researchers in this community to build upon our work and advance the field.